OpenStudy (anonymous):

Simplify the expression 20 - 4 · 2 ___________ 12 divided by 4 + 1 A. 2 B. 3 C. 8 D. 4

OpenStudy (ashleyisakitty):

First do the order of operations on each equation.

OpenStudy (igreen):

Is it this? \(\dfrac{24 - 4 * 2}{12 \div 4 + 1}\)

OpenStudy (ashleyisakitty):

P.E.M.D.A.S. Parenthesis, Exponents, Multiplication/Division, then addition/subtraction.

OpenStudy (ashleyisakitty):

@JillyCraft64 have you done the first step?

OpenStudy (igreen):

\(\dfrac{24 - 4 * 2}{12 \div 4 + 1}\) First multiply 4 and 2 in the numerator: 4 * 2 = 8 \(\dfrac{24 - 8}{12 \div 4 + 1}\) Now divide 12 and 4 in the denominator: 12 / 4 = 3 \(\dfrac{24 - 8}{3 + 1}\) Now subtract 24 and 8 in the numerator: 24 - 8 = 16 \(\dfrac{16}{3 + 1}\) Add 3 and 1 in the denominator: 3 + 1 = 4 \(\dfrac{16}{4}\) Now divide 16 and 4: 16 / 4 = 4 \(4\)

OpenStudy (ashleyisakitty):

-.- okay or just give away the answer process... thats cool too.

OpenStudy (igreen):

Lol..I didn't see your responses because my text become so long..sorry. :P

OpenStudy (igreen):

However I had already explained PEMDAS to him on his previous question. @Ashleyisakitty

OpenStudy (ashleyisakitty):

if so, you need to allow the user to work on the process after explanation, or else they wont learn.